In x minutes, there will be f(x) bacteria present in a certain culture, where f(x)=Ae raised to 0.03x. If 100,000 bacteria are present initially,
a) How many will be present in 1 minute?
b) After how many minutes will there be 200,000 present?

Answer by Alan3354(69443)

In x minutes, there will be f(x) bacteria present in a certain culture, where f(x)=Ae raised to 0.03x. If 100,000 bacteria are present initially,
a) How many will be present in 1 minute?
b) After how many minutes will there be 200,000 present?
---------------------
f(x)=Ae raised to 0.03x
Use math terms:
f(x)=Ae^(0.03x) = 100000*e^(0.03x)
---
a) How many will be present in 1 minute?
Sub 1 for x.
================================
b) After how many minutes will there be 200,000 present?
f(x)=100000e^(0.03x) = 200000
Solve for x